OpenStudy (anonymous):

HELLLPP!! What is the equation of the line that passes through (2, 1) and is perpendicular to the graph of y = –x + 2?

OpenStudy (anonymous):

The slope would be 1 of that line as we know if m1 is a slope of a starighline1 and m2 is a straightline2 and if they are perpendicular we have m1*m2=-1....so thats how we get the slope...now we know that for any equation of starightline we can write y=mx + c ,here m=1 now put the points and get c...and ya will get the straightline

OpenStudy (anonymous):

can youu help me out ?? im stuck! ;?

OpenStudy (anonymous):

@arghya

OpenStudy (anonymous):

y=x - 1

OpenStudy (anonymous):

thnx! ;)